How many liters of a 60% alcohol solution must be mixed with 50 liters of a 90% solution to get a 70% solution?
Answer by stanbon(75887)

How many liters of a 60% alcohol solution must be mixed with 50 liters of a 90% solution to get a 70% solution?
--------------------------------------
Equation:
alcohol+ alcohol = alcohol
0.60x + 0.90*50 = 0.70(x+50)
Multiply thru by 100 to get:
60x + 90*50 = 70x + 70*50
------
-10x = -20*50
x = 100 liters (amt. of 60% solution needed in the mixture)
===========================
Cheers,
Stan H.